OpenStudy (anonymous):

Joint d.f of X and Y: F(x,y)=(1/156) xy(x^2+y) ,0<=x<=3,0<=y<=4 determine distribution function of Y

OpenStudy (anonymous):

F(x,y)=(1/156) xy(\[x^{2}\]+y) ;\[0\le x \le3,0\le y \le4\]

OpenStudy (anonymous):

Have a look here -> http://en.wikipedia.org/wiki/Cumulative_distribution_function